OpenStudy (anonymous):

Find three solutions of the equation. y = –2x (0, 0), (3, –8), (–2, –4) (1, –2), (0, 0), (–1, 1) (–2, 4), (1, –2), (0, 0) (–2, 4), (1, –2), (2, –3) Find the slope of the line through (3, –6) and (5, 8). –7 7 - 1/7 1/7

OpenStudy (anonymous):

(–2, 4), (1, –2), (0, 0) You just plug in numbers. Part II: Use the slope formula. 8 + 6/5 - 3 = 14/2 = 7 Slope = 7

OpenStudy (anonymous):

Thank you for your help. :)

OpenStudy (anonymous):

You're welcome :)